Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
At what time will John catch up with Martina?

Answer by checkley71(8403)

DISTANCE=RATE*TIME
SEEING AS THESE DISTANCES ARE THE SAME WE SET UP THE EQUATION:
24T=48(T-2)
24T=48T-96
-24T=-96
T=96/24
T=4 HOURS AFTER NOON THEY WILL MEET.
PROOF
24*4=48(4-2)
96=48*2
96=96